You have two vertical angles. One measures (4x+2) degrees. The other measures (5x-13) degrees. What is the measure of each angle?
a.) 62 and 62
b.) 28 and 28
c.) 118 and 118
d.) 19 and 19

Answer by dolly(163)

Vertical angles are always equal
==> 5x - 13 = 4x + 2
==> 5x - 4x = 2 + 13
==> x = 15
So each angle = 4x + 2
= 4(15) + 2
= 62
Thus each angle is 62, 62
good luck